OpenStudy (anonymous):

yx^3/(x^-3y^0 x x^-2y^2)^-2 Simplify your answer should contain only positive exponents.

OpenStudy (anonymous):

I'm assuming that this the correct form of the expression: \[yx^3/(x^{-3}.y^0.x^{-2}.y^2)^{-2}\] anything to the power 0 is 1, so that y^0 just disappears, combine the powers of x in the denominator to get x^(-5) now apply the power of power in the whole denominator to get \[yx^3/(x^{-5}.y^2)^{-2} = yx^3/(x^{10}.y^{-4})\] now deal with combining the powers of x and powers of y from the top & bottom \[y^{1--4}.x^{3-10} = y^5.x^{-7} = y^5/x^7\]